# Moduli stacks of Higgs bundles on stable curves

### v2

@@ 8.1.2 Ordinary Artin Stack of Gieseker-Higgs bundles
$M^{cl}_{Gie}\longrightarrow TFH(\mathcal{X}/S)$ (8.2)
$(\mathfrak{X},\mathcal{E},\phi)\mapsto(\pi_{*}\mathcal{E},\pi_{*}\phi)$ (8.3)
@@
8.1.3 Construction of an atlas for $N_{Gie}$
Now, we recall the construction of an atlas for the stack $M^{{}^{cl}}_{Gie}$ as well as the classical stack of Gieseker vector bundles. Let us denote by $N_{Gie}$ the classical stack of Gieseker vector bundles.
Let $\mathcal{O}_{\mathcal{X}/S}(1)$ be a relatively ample line bundle for the family of curves $\mathcal{X}/S$. The set of all flat families of stable torsion-free sheaves (Higgs pairs) of degree $d$ and rank $n$ over $\mathcal{X}$ forms a bounded family. Therefore we can choose a large integer $m_{0}$ such that given any family of stable torsion-free Higgs pairs $(\mathcal{F}_{S},\phi_{S})$, the sheaf $\mathcal{F}_{s}\otimes\mathcal{O}_{\mathcal{X}_{s}}(m_{0})$ is generated by global sections and $H^{1}(\mathcal{X}_{s},\mathcal{F}_{s}\otimes\mathcal{O}_{\mathcal{X}_{s}}(m_{0}))=0$ for every geometric point $s\in S$. Set $N(m_{0}):=H^{0}(\mathcal{X}_{s},\mathcal{F}_{s}\otimes\mathcal{O}_{\mathcal{X}_{s}}(m_{0}))$ for any geometric point $s\in S$. We denote by $Grass(N(m_{0}),n)$ the Grassmannian of $n$ dimensional quotient vector spaces of $\mathbb{k}^{N(m_{0})}$.
Definition 8.6.
Let $\mathcal{G}^{m_{0}}_{S}:{Sch}/S\longrightarrow{Sets}$ be the functor defined as follows:
$\mathcal{G}^{m_{0}}_{S}(T)=\{(\Delta_{T},V_{T})\},$ (8.4)
@@ 8.1.3 Construction of an atlas for $N_{Gie}$
$\Delta_{T}\subset\mathcal{X}\times_{S}T\times{Grass}(N(m_{0}),n)$ (8.5)
is a closed subscheme, and $V_{T}$ is a vector bundle on $\Delta_{T}$ such that
  1. 1.
    the projection $j:\Delta_{T}\longrightarrow T\times{Grass}(N(m_{0}),n)$ is a closed immersion,
  2. 2.
    the projection $\Delta_{T}\longrightarrow\mathcal{X}\times_{S}T$ is a modification,
  3. 3.
    the projection $p_{T}:\Delta_{T}\longrightarrow T$ is a flat family of Gieseker curves,
  4. 4.
@@
    Let $\mathcal{V}$ be the tautological quotient bundle of rank $n$ on ${Grass}(N(m_{0}),n)$ and $\mathcal{V}_{T}$ its pullback to $T\times{Grass}(N(m_{0}),n)$. Then
    $V_{T}:=j^{*}(\mathcal{V}_{T})$ (8.6)
    be such that $V_{T}$ is a Gieseker vector bundle on the modification $\Delta_{T}$ of rank $n$ and degree $d(m_{0})^{\prime}:=N(m_{0})+n(g-1)$.
  5. 5.
    for each $t\in T$, the quotient $\mathcal{O}_{\Delta_{t}}^{N}(m_{0})\longrightarrow V_{t}$ induces an isomorphism
    $H^{0}(\Delta_{t},\mathcal{O}_{\Delta_{t}}^{N(m_{0})})\cong H^{0}(\Delta_{t},V_{t})$ (8.7)
    and $H^{1}(\Delta_{t},V_{t})=0$.
@@
We denote by $P(m_{0})$ the Hilbert polynomial of the closed subscheme $\Delta_{s}$ of $\mathcal{X}_{s}\times Grass(N(m_{0}),n)$ for any geometric point $s\in S$ with respect to the polarisation $\mathcal{O}_{\mathcal{X}_{s}}(1)\boxtimes\mathcal{O}_{Grass(N(m_{0}),n)}(1)$, where $\mathcal{O}_{Grass(N(m_{0}),n)}(1)$ is the line bundle $\det\leavevmode\nobreak\ \leavevmode\nobreak\ \mathcal{V}$.
@@
It is shown in [25, Proposition 8] that the functor $\mathcal{G}^{m_{0}}_{S}$ is represented by a $PGL(N(m_{0}))$-invariant open subscheme $\mathcal{Y}^{m_{0}}_{S}$ of the Hilbert scheme $\mathcal{H}_{S}:={Hilb}^{P(m_{0})}(\mathcal{X}\times{Grass}(N(m_{0}),n))$.
Finally, the disjoint union $\mathcal{Y}_{S}:=(\underset{N(m),m\geq m_{0}}{\coprod}\mathcal{Y}^{m}_{S})\longrightarrow N_{Gie}$ is an atlas. The varieties $\mathcal{Y}^{m}_{S}$ are all smooth varieties and the projection map to the stack is also smooth.
8.1.4 Construction of an atlas for $M^{{}^{cl}}_{Gie}$
Definition 8.7.
We define a functor
$\mathcal{G}^{H,m}_{S}:{Sch}/\mathcal{Y}_{S}\longrightarrow{Groups}$ (8.8)
which maps
$T\longrightarrow H^{0}(T,(p_{T})_{*}(\mathcal{E}nd\leavevmode\nobreak\ \leavevmode\nobreak\ \mathcal{V}_{T}\otimes\omega_{\Delta_{T}/T})),$
@@ 8.1.4 Construction of an atlas for $M^{{}^{cl}}_{Gie}$
where $p_{T}:\Delta_{T}:=\Delta_{{}_{\mathcal{Y}_{S}}}\times_{{}_{\mathcal{Y}_{S}}}T\longrightarrow T$ is the projection, and $\omega_{\Delta_{T}/T}$ denotes the relative dualising sheaf of the family of curves $p_{T}$.
@@
Since $\mathcal{Y}^{m}_{S}$ is a reduced scheme, the functor $\mathcal{G}^{H,m}_{S}$ is representable, i.e., there exists a linear $\mathcal{Y}^{m}_{S}$- scheme $\mathcal{Y}^{H,m}_{S}$, which represents it. For a $S$-scheme $T$, a point in $\mathcal{G}^{H,m}_{S}(T)$ is given by $(V_{T},\phi_{T})$, where
  1. 1.
    $V_{T}\in\mathcal{G}^{m}_{S}(T)$, and
  2. 2.
    $(V_{T},\phi_{T})$ is a Gieseker-Higgs bundle.
Finally, the disjoint union $\mathcal{Y}^{H}_{S}:=(\underset{N(m),m\geq m_{0}}{\coprod}\mathcal{Y}^{H,m}_{S})\longrightarrow M^{{}^{cl}}_{Gie}$ is an atlas. The projection map to the stack is smooth.
Remark 8.8.
Notice that there is a morphism $\tilde{\theta}:\mathcal{Y}^{H,m}_{S}\to R^{\Lambda,m}_{S}$ given by pushforward
$\tilde{\theta}:\mathcal{Y}^{H,m}_{S}\to R^{\Lambda,m}_{S}$ (8.9)
$(\Delta_{T},V_{T},\phi_{T})\mapsto((\pi_{T})_{*}V_{T},(\pi_{T})_{*}\phi_{T})$ (8.10)
@@
where $\pi_{T}:\Delta_{T}\longrightarrow\mathcal{X}\times_{S}T$ is the modification map and $T$ is an $S$-scheme. It is straightforward to check that the map is $GL_{N(m)}$-equivariant morphism.
8.2 Dimension and local properties of $M^{cl}_{Gie,0}$
@@
8.2.1 Relative log-symplectic reduction
@@
We will do this by a log-symplectic reduction on every atlas of $N_{Gie,0}$. Let us pick one of the atlas $\mathcal{Y}^{m}_{S}$ of $M^{cl}_{Gie}$. The group $GL_{N(m)}$ acts on $\mathcal{Y}^{m}_{S}$ and the quotient stack $[\mathcal{Y}^{m}_{S}/GL_{N(m)}]$ is an open sub-stack of $N_{Gie}$. We consider the relative log-cotangent bundle $\Omega^{\log}_{\mathcal{Y}^{m}_{S}/S}$. Since the $GL_{N(m)}$ action preserves the normal crossing divisor $\mathcal{Y}^{m}_{0}$ (the closed fibre of $\mathcal{Y}^{m}_{S}\longrightarrow S$), the action of $GL_{N(m)}$ lifts to an action on $\Omega^{\log}_{\mathcal{Y}^{m}_{S}/S}$ with a moment map $\mu_{\log}:\Omega^{\log}_{\mathcal{Y}^{m}_{S}/S}\longrightarrow\mathfrak{gl}_{N(m)}^{*}$. Notice that the action of $GL_{N(m)}$ has a generic stabiliser $\mathbb{G}_{m}$; therefore the map $\mu_{\log}$ actually factors through $\Omega^{\log}_{\mathcal{Y}^{m}_{S}/S}\longrightarrow\mathfrak{pgl}_{N(m)}^{*}$. Since we know that there exists an open subset of $\mathcal{Y}^{m}_{S}$ where the action of $GL_{N(m)}$ has stabiliser isomorphic to $\mathbb{G}_{m}$ (namely, the locus of stable vector bundles), therefore the map $\Omega^{log}_{\mathcal{Y}^{m}_{S}/S}\longrightarrow\mathfrak{pgl}_{N(m)}^{*}$ is surjective. Then one can show that $\mathcal{Y}^{H,m}_{S}=\mu_{\log}^{-1}(0)$. To see this, notice that $\mathcal{Y}^{m}_{S}$ is a principal $GL_{N(m)}$-bundle over an open subset of the stack $N_{Gie}$. Let $a$ denote the projection $a:\mathcal{Y}^{m}_{S}\longrightarrow N_{Gie}$. Let us write the cotangent sequence.
$0\longrightarrow a^{*}\Omega^{\log,*}_{N_{Gie}/S}\longrightarrow T^{\log,*}_{\mathcal{Y}^{m}_{S}/S}\longrightarrow\mathfrak{g}l^{*}_{N(m)}\otimes\mathcal{O}_{\mathcal{Y}^{m}_{S}}$ (8.11)
@@
where the last map is given by $\mu_{\log}$. Notice that $[T^{\log,*}_{\mathcal{Y}^{m}_{S}/S}\longrightarrow\mathfrak{g}l^{*}_{N(m)}\otimes\mathcal{O}_{\mathcal{Y}^{m}_{S}}]$ is the cotangent complex of the stack $N_{Gie}$. Therefore, using standard arguments from deformation theory of vector bundles it follows that the 00-th cohomology of the complex at a point $[\mathcal{O}^{N(m)}\rightarrow\mathcal{E}]\in\mathcal{Y}^{m}_{S}$ is isomorphic to $H^{1}(\mathcal{E}nd\mathcal{E})^{\vee}\cong\mathsf{Hom}(\mathcal{E},\mathcal{E}\otimes\omega)$. Therefore by Definition 8.7, $\mathcal{Y}^{H,m}_{S}=\mu_{\log}^{-1}(0)$. Therefore $[\mu_{\log}^{-1}(0)/GL_{N(m)}]$ is an open subset of $M^{cl}_{Gie}$. Therefore, we will compute the dimension of $\mu_{\log}^{-1}(0)$ and that it is a local complete intersection. We recall a result on the dimension of the image of a cotangent fibre under the moment map.
Lemma 8.9.
@@
$\dim\mu_{\log}(\Omega^{\log}_{\mathcal{Y}^{m}_{S}/S,y})\geq\leavevmode\nobreak\ \dim\leavevmode\nobreak\ \big{(}\frac{\mathfrak{gl}_{N(m)}}{\mathfrak{gl}_{N(m),y}}\big{)}^{*}$, where $\mathfrak{gl}_{N(m),y}$ is the Lie algebra of the stabilizer of $y\in\mathcal{Y}^{m}_{S}$ under the action of $GL_{N}(m)$.
@@
Here $K$ denotes the kernel of the natural map $T^{log}_{\mathcal{Y}^{m}_{S}/S,y}\longrightarrow T_{\mathcal{Y}^{m}_{S}/S,y}$. Notice that the map $\mathfrak{gl}_{N(m)}\longrightarrow T_{\mathcal{Y}^{m}_{S}/S,y}$ is the differential of the orbit map, and it factors through $T^{log}_{\mathcal{Y}^{m}_{S}/S,y}$ because the action of $GL_{N(m)}$ preserves the normal crossing divisor. It is well known that the rightmost vertical map is injective [37, Lemma 2.4.1]. Therefore we can complete the diagram.
@@
Therefore we see that $\ker(\iota_{log})\subset\ker(\iota)=\mathfrak{gl}_{N(m),y}$. Since the moment and log-moment maps are the dual of the maps $\iota$ and $\iota_{log}$, therefore $\dim\leavevmode\nobreak\ \mu_{log}(\Omega^{log}_{\mathcal{Y}^{m}_{S}/S,y})\geq\leavevmode\nobreak\ \dim\leavevmode\nobreak\ \big{(}\frac{\mathfrak{gl}_{N(m)}}{\mathfrak{gl}_{N(m),y}}\big{)}^{*}$. ∎
Proposition 8.10.
The closed fibre $N^{cl}_{Gie,0}$ is an irreducible, equidimensional, almost very good stack ([37, Definition 2.1.2]) with normal crossing singularities.
@@
From [19, Theorem 9.5],it follows that the normalisation of $N^{cl}_{Gie,0}$ is a $KGL_{n}$ bundle over the stack of vector bundles $\mathrm{Bun}(\tilde{X}_{0})$ of rank $n$ and degree $d$ over the normalisation $\tilde{X}_{0}$ of the nodal curve $X_{0}$. Here $KGL_{n}$ denotes the compactification of $GL_{n}$ constructed by Kausz [20]. More precisely, let $E$ is the universal $GL_{n}$ bundle over $\tilde{X_{0}}\times\mathrm{Bun}(\tilde{X_{0}})$. Consider the $GL_{n}\times GL_{n}$ bundle $E_{x_{1}}\times_{\mathrm{Bun}(\tilde{X_{0}})}E_{x_{2}}$ over $\mathrm{Bun}(\tilde{X_{0}})$. Then the associated $KGL_{n}$ fibration $(E_{x_{1}}\times_{Bun(\tilde{X_{0}})}E_{x_{2}})\times_{GL_{n}\times GL_{n}}KGL_{n}\cong\widetilde{N^{cl}_{Gie,0}}$. This is the stack of Gieseker vector bundle data ([19, Definition 4.7]), which is equivalent to ([10, Lemma 5.6]) the stack of marked Gieseker vector bundles (i.e., a Gieseker vector bundle with a marked node ([10, Definition 5.1])). It is obvious that the automorphism group of a marked Gieseker vector bundle is isomorphic to the automorphism group of the corresponding Gieseker vector bundle. Therefore, it is enough to show that the stack of Gieseker vector bundle data is smooth, equidimensional, irreducible and almost very good.
Now let us recall that the map $\tilde{\pi}:\widetilde{N^{cl}_{Gie,0}}\longrightarrow\mathrm{Bun}(\tilde{X}_{0})$ is given by $(X^{m,n},s_{1},s_{2},\tilde{\mathcal{E}},\phi)\mapsto h_{*}(\mathcal{E}(-s_{1}-s_{2}))(p_{1}+p_{2})$, where $h:X^{m,n}\longrightarrow\tilde{X}_{0}$ is the modification map ([19, Lemma 9.3]). Here $(X^{m,n},s_{1},s_{2},\tilde{\mathcal{E}},\phi)$ is a Gieseker vector bundle data and here $\phi$ is not a Higgs bundle but an identification between the fibres ~$\tilde{\mathcal{E}}_{s_{1}}\xrightarrow{\cong}\tilde{\mathcal{E}}_{s_{2}}$ ([19, Definition 4.7]). It is straightforward to check that the induced map
@@
is injective. For notational convenience, let us denote $Y:=\widetilde{N^{cl}_{Gie,0}}$ and $Z:=\mathrm{Bun}(\tilde{X}_{0})$.
From [4, Proposition 2.1.2] and [37, Definition 2.1.2, Remark 2.1.3], it follows that $Z$ is smooth, irreducible, almost very good stack. More precisely, the $\operatorname{codim}_{Z}(Z_{k})>k\leavevmode\nobreak\ \leavevmode\nobreak\ \forall k\geq 1$, where $Z_{k}:=\{z\in Z\leavevmode\nobreak\ |\leavevmode\nobreak\ \text{dim}\leavevmode\nobreak\ \operatorname{Aut}\leavevmode\nobreak\ (z)=1+k\}$. From the fact that $\operatorname{Aut}(X^{m,n},s_{1},s_{2},\tilde{\mathcal{E}},\phi)\subset\operatorname{Aut}(h_{*}(\mathcal{E}(-s_{1}-s_{2}))(p_{1}+p_{2}))$, it follows that $\tilde{\pi}(Y_{k})\subset Z_{k}$. Therefore, $Y_{k}\subset\tilde{\pi}^{-1}(Z_{k})$ and
$\dim\leavevmode\nobreak\ Y_{k}\leq\dim\leavevmode\nobreak\ \leavevmode\nobreak\ \tilde{\pi}^{-1}(Z_{k})=\dim\leavevmode\nobreak\ \leavevmode\nobreak\ Z_{k}+\dim\leavevmode\nobreak\ \leavevmode\nobreak\ KGL_{n}.$
@@
$\begin{split}\operatorname{codim}_{Y}(Y_{k})=&\dim\leavevmode\nobreak\ \leavevmode\nobreak\ Y-\dim\leavevmode\nobreak\ \leavevmode\nobreak\ Y_{k}\\ &\geq\leavevmode\nobreak\ \leavevmode\nobreak\ \dim\leavevmode\nobreak\ \leavevmode\nobreak\ Y-\dim\leavevmode\nobreak\ \leavevmode\nobreak\ Z_{k}-\dim\leavevmode\nobreak\ \leavevmode\nobreak\ KGL_{n}=\dim\leavevmode\nobreak\ \leavevmode\nobreak\ Z-\dim\leavevmode\nobreak\ \leavevmode\nobreak\ Z_{k}\\ &=\operatorname{codim}_{Z}(Z_{k})>k.\end{split}$
Therefore, $\widetilde{N^{cl}_{Gie,0}}$ is smooth irreducible, almost very good stack and $N^{cl}_{Gie,0}$ is an irreducible, almost very good stack. ∎
Theorem 8.11.
The stack $M^{cl}_{Gie,0}$ is an irreducible local complete intersection of pure dimension $12\cdot\text{dim}\leavevmode\nobreak\ \leavevmode\nobreak\ N_{Gie,0}+$ 1.

### v3

@@ A.1.3. Classical Artin stack of Gieseker-Higgs bundles
(A.2) $M^{cl}_{Gie}\longrightarrow\mathsf{TFH}(\mathcal{X}/S)$
(A.3) $(\mathfrak{X},\mathcal{E},\phi)\mapsto(\pi_{*}\mathcal{E},\pi_{*}\phi)$
@@
Remark A.6.
We remind the reader that the stacks $M^{cl}$ and $M^{cl}_{Gie}$ are defined over the stack of expanded degenerations $\mathfrak{M}$. But the codomain stacks $\operatorname{\mathsf{Coh}}(\mathcal{X}/S)$ and $\mathsf{TFH}(\mathcal{X}/S)$ live only on $S$.
A.2. Construction of an atlas for $N_{Gie}$
Let us denote by $N_{Gie}$ the classical stack of Gieseker vector bundles. It can be seen as the closed substack of $M^{cl}_{Gie}$ consisting of the Gieseker-Higgs bundles $(\mathcal{E},\phi)$, whose Higgs field $\phi=0$. In this subsection, we will give the construction of an atlas of the stack $N_{Gie}$. Let $\mathcal{O}_{\mathcal{X}/S}(1)$ be a relatively ample line bundle for the family of curves $\mathcal{X}/S$. As before, we fix rank $n$ and degree $d$.
Definition A.7.
Let $m,N(m)$ be positive integers with $N(m)\geq n$. Define the functor
$\mathcal{G}^{m}_{S}:{Sch}/S\longrightarrow{Sets}$
by
$\mathcal{G}^{m}_{S}(T)=\{(\Delta_{T},V_{T})\},$
@@ A.2. Construction of an atlas for $N_{Gie}$
    $\Delta_{T}\subset\mathcal{X}\times_{S}T\times\mathrm{Grass}(N(m),n)$ is a closed subscheme,
  2. (2)
    the projection $j:\Delta_{T}\to T\times\mathrm{Grass}(N(m),n)$ is a closed immersion,
  3. (3)
    the projection $\Delta_{T}\to\mathcal{X}\times_{S}T$ is a modification,
  4. (4)
    the projection $p_{T}:\Delta_{T}\to T$ is a flat family of Gieseker curves,
  5. (5)
@@
    Let $\mathcal{V}$ be the tautological quotient bundle of rank $n$ on $\mathrm{Grass}(N(m),n)$ and $\mathcal{V}_{T}$ its pullback to $T\times\mathrm{Grass}(N(m),n)$. Then
    $V_{T}:=j^{*}(\mathcal{V}_{T})(-m)$
    is a Gieseker vector bundle on the modification $\Delta_{T}$ of rank $n$ and degree $d$.
  6. (6)
    For each $t\in T$, the quotient $\mathcal{O}_{\Delta_{t}}^{N(m)}\twoheadrightarrow V_{t}(m)$ induces an isomorphism
    $H^{0}(\Delta_{t},\mathcal{O}_{\Delta_{t}}^{N(m)})\;\cong\;H^{0}(\Delta_{t},V_{t}(m))$
    and $H^{1}(\Delta_{t},V_{t}(m))=0$.
@@
Let $P(m)$ be the Hilbert polynomial of the closed subscheme $\Delta_{s}$ of $\mathcal{X}_{s}\times\mathrm{Grass}(N(m),n)$ with respect to the polarisation $\mathcal{O}_{\mathcal{X}_{s}}(1)\boxtimes\mathcal{O}_{\mathrm{Grass}(N(m),n)}(1)$, where $\mathcal{O}_{\mathrm{Grass}(N(m),n)}(1)=\det\mathcal{V}$.
@@
It is shown in [25, Proposition 8] that the functor $\mathcal{G}^{m}_{S}$ is represented by a $PGL(N(m))$-invariant open subscheme $\mathcal{Y}^{m}_{S}$ of the Hilbert scheme
$\mathcal{H}_{S}=\mathrm{Hilb}^{P(m)}(\mathcal{X}\times\mathrm{Grass}(N(m),n)).$
Finally, the disjoint union
$\mathcal{Y}_{S}:=\coprod_{N(m),\,m\geq 0}\mathcal{Y}^{m}_{S}\longrightarrow N_{Gie}$
is an atlas. The varieties $\mathcal{Y}^{m}_{S}$ are smooth and the projection map to the stack is smooth.
A.3. Construction of an atlas for $M^{cl}_{Gie}$
Definition A.8.
Define the functor
$\mathcal{G}^{H,m}_{S}:{Sch}/\mathcal{Y}_{S}\longrightarrow{Groups}$
by
$\mathcal{G}^{H,m}_{S}(T)=H^{0}\bigl(T,(p_{T})_{*}(\mathcal{E}\mathsf{nd}\mathcal{V}_{T}\otimes\omega_{\Delta_{T}/T})\bigr),$
@@ A.3. Construction of an atlas for $M^{cl}_{Gie}$
where $p_{T}:\Delta_{T}:=\Delta_{\mathcal{Y}_{S}}\times_{\mathcal{Y}_{S}}T\to T$ is the projection, and $\omega_{\Delta_{T}/T}$ is the relative dualizing sheaf of the family of curves $p_{T}$.
@@
Since each $\mathcal{Y}^{m}_{S}$ is reduced, the functor $\mathcal{G}^{H,m}_{S}$ is representable: there exists a linear scheme $\mathcal{Y}^{H,m}_{S}$ over $\mathcal{Y}^{m}_{S}$ representing it.
For an $S$-scheme $T$, a point in $\mathcal{G}^{H,m}_{S}(T)$ consists of $(V_{T},\phi_{T})$ where $(V_{T})$ is in $\mathcal{G}^{m}_{S}(T)$ and $(V_{T},\phi_{T})$ is a Gieseker–Higgs bundle.
The disjoint union
$\mathcal{Y}^{H}_{S}:=\coprod_{N(m),\,m\geq 0}\mathcal{Y}^{H,m}_{S}\longrightarrow M^{cl}_{Gie}$
is an atlas. The projection map to the stack is smooth.
Remark A.9.
There is a natural morphism
$\tilde{\theta}:\mathcal{Y}^{H,m}_{S}\longrightarrow R^{\Lambda,m}_{S}$
given by pushforward
$(\Delta_{T},V_{T},\phi_{T})\;\longmapsto\;\bigl((\pi_{T})_{*}V_{T},\;(\pi_{T})_{*}\phi_{T}\bigr),$
@@
where $\pi_{T}:\Delta_{T}\to\mathcal{X}\times_{S}T$ is the modification map. This morphism is $GL_{N(m)}$-equivariant.
A.4. Dimension and local properties of $M^{cl}_{Gie,0}$
@@
A.4.1. Relative log-symplectic reduction
Fix one atlas component $\mathcal{Y}^{m}_{S}$ of $M^{cl}_{Gie}$. The group $GL_{N(m)}$ acts on $\mathcal{Y}^{m}_{S}$ and the quotient stack $\mathcal{Y}^{m}_{S}/GL_{N(m)}]$ is an open sub-stack of $N_{Gie}$.
@@ A.4.1. Relative log-symplectic reduction
Consider the relative log-cotangent bundle $\Omega^{\log}_{\mathcal{Y}^{m}_{S}/S}$ of the morphism $\mathcal{Y}^{m}_{S}\to S$. Since the $GL_{N(m)}$ action preserves the normal crossing divisor $\mathcal{Y}^{m}_{0}$ (the closed fibre of $\mathcal{Y}^{m}_{S}\longrightarrow S$), the action of $GL_{N(m)}$ lifts to an action on $\Omega^{\log}_{\mathcal{Y}^{m}_{S}/S}$ with a moment map $\mu_{\log}:\Omega^{\log}_{\mathcal{Y}^{m}_{S}/S}\longrightarrow\mathfrak{gl}_{N(m)}^{*}$. The action of $GL_{N(m)}$ has a generic stabiliser $\mathbb{G}_{m}$; therefore the map $\mu_{\log}$ actually factors through $\Omega^{\log}_{\mathcal{Y}^{m}_{S}/S}\longrightarrow\mathfrak{pgl}_{N(m)}^{*}$. Since we know that there exists an open subset of $\mathcal{Y}^{m}_{S}$ where the action of $GL_{N(m)}$ has stabiliser isomorphic to $\mathbb{G}_{m}$ (namely, the locus of stable vector bundles), therefore the map $\Omega^{log}_{\mathcal{Y}^{m}_{S}/S}\longrightarrow\mathfrak{pgl}_{N(m)}^{*}$ is surjective. Then one can show that
(A.4) $\mathcal{Y}^{H,m}_{S}=\mu_{\log}^{-1}(0).$
To see this, note that $\mathcal{Y}^{m}_{S}$ is a principal $GL_{N(m)}$ bundle over an open subset of the stack $N_{Gie}$. Let $a$ denote the projection $a:\mathcal{Y}^{m}_{S}\longrightarrow N_{Gie}$. Let us write the cotangent sequence of the sheaf of relative logarithmic differential forms.
(A.5) $0\longrightarrow a^{*}(T^{\log}_{N_{Gie}/S})^{*}\longrightarrow(T^{\log}_{\mathcal{Y}^{m}_{S}/S})^{*}\longrightarrow\mathfrak{g}l^{*}_{N(m)}\otimes\mathcal{O}_{\mathcal{Y}^{m}_{S}}$
@@
where the last map is given by $\mu_{\log}$. Notice that $(T^{\log}_{\mathcal{Y}^{m}_{S}/S})^{*}\longrightarrow\mathfrak{g}l^{*}_{N(m)}\otimes\mathcal{O}_{\mathcal{Y}^{m}_{S}}]$ is the cotangent complex of the stack $N_{Gie}$. Therefore, using standard arguments from the deformation theory of vector bundles, it follows that the 0-th cohomology of the complex at a point $\mathcal{O}^{N(m)}\rightarrow\mathcal{E}]\in\mathcal{Y}^{m}_{S}$ is isomorphic to $H^{1}(\mathcal{E}\mathsf{nd}\mathcal{E})^{\vee}\cong\mathsf{Hom}(\mathcal{E},\mathcal{E}\otimes\omega)$. Therefore, by Definition A.8, $\mathcal{Y}^{H,m}_{S}=\mu_{\log}^{-1}(0)$.
Therefore,
$\mu_{\log}^{-1}(0)/GL_{N(m)}]$
is an open subset of $M^{cl}_{Gie}$.
We will compute the dimension of $\mu_{\log}^{-1}(0)$ and show that it is local complete intersection. We begin by recalling a result on the dimension of the image of a cotangent fibre under the moment map.
Lemma A.10.
@@
$\dim\mu_{\log}(\Omega^{\log}_{\mathcal{Y}^{m}_{S}/S,y})\geq\penalty 10000\ \dim\penalty 10000\ \big(\frac{\mathfrak{gl}_{N(m)}}{\mathfrak{gl}_{N(m),y}}\big)^{*}$, where $\mathfrak{gl}_{N(m),y}$ is the Lie algebra of the stabilizer of $y\in\mathcal{Y}^{m}_{S}$ under the action of $GL_{N}(m)$.
@@
Here $K$ denotes the kernel of the natural map $T^{\log}_{\mathcal{Y}^{m}_{S}/S,y}\longrightarrow T_{\mathcal{Y}^{m}_{S}/S,y}$. Notice that the map $\mathfrak{gl}_{N(m)}\longrightarrow T_{\mathcal{Y}^{m}_{S}/S,y}$ is the differential of the orbit map, and it factors through $T^{log}_{\mathcal{Y}^{m}_{S}/S,y}$ because the action of $GL_{N(m)}$ preserves the normal crossing divisor. It is well known that the rightmost vertical map is injective [39, Lemma 2.4.1]. Therefore, we can complete the diagram.
@@
Therefore we see that $\ker(\iota_{\log})\subset\ker(\iota)=\mathfrak{gl}_{N(m),y}$. Since moment and log-moment maps are the dual of the maps $\iota$ and $\iota_{\log}$, therefore $\dim\penalty 10000\ \mu_{\log}(\Omega^{\log}_{\mathcal{Y}^{m}_{S}/S,y})\geq\penalty 10000\ \dim\penalty 10000\ \big(\frac{\mathfrak{gl}_{N(m)}}{\mathfrak{gl}_{N(m),y}}\big)^{*}$. ∎
Proposition A.11.
The closed fibre $N^{cl}_{Gie,0}$ is an irreducible, equidimensional, almost very good stack ([39, Definition 2.1.2]) with normal crossing singularities.
@@
[…]
Now let us recall that the map $\tilde{\pi}:\widetilde{N^{cl}_{Gie,0}}\longrightarrow\mathrm{Bun}(\tilde{X}_{0})$ is given by $(X^{m,n},s_{1},s_{2},\tilde{\mathcal{E}},\phi)\mapsto h_{*}(\mathcal{E}(-s_{1}-s_{2}))(p_{1}+p_{2})$, where $h:X^{m,n}\longrightarrow\tilde{X}_{0}$ is the modification map ([18, Lemma 9.3]). Here $(X^{m,n},s_{1},s_{2},\tilde{\mathcal{E}},\phi)$ is a Gieseker vector bundle data and here $\phi$ is not a Higgs bundle but an identification between the fibres $\tilde{\mathcal{E}}_{s_{1}}\xrightarrow{\cong}\tilde{\mathcal{E}}_{s_{2}}$ ([18, Definition 4.7]). It is straightforward to check that the induced map
@@
is injective. For notational convenience, denote $Y:=\widetilde{N^{cl}_{Gie,0}}$ and $Z:=\mathrm{Bun}(\tilde{X}_{0})$.
From [5, Proposition 2.1.2] and [39, Definition 2.1.2, Remark 2.1.3], it follows that $Z$ is smooth, irreducible, almost very good stack. More precisely, the $\operatorname{codim}_{Z}(Z_{k})>k\penalty 10000\ \penalty 10000\ \forall k\geq 1$, where $Z_{k}:=\{z\in Z\penalty 10000\ |\penalty 10000\ \dim\penalty 10000\ \operatorname{Aut}\penalty 10000\ (z)=1+k\}$. From the fact that $\operatorname{Aut}(X^{m,n},s_{1},s_{2},\tilde{\mathcal{E}},\phi)\subset\operatorname{Aut}(h_{*}(\mathcal{E}(-s_{1}-s_{2}))(p_{1}+p_{2}))$, it follows that $\tilde{\pi}(Y_{k})\subset Z_{k}$. Therefore, $Y_{k}\subset\tilde{\pi}^{-1}(Z_{k})$ and
$\dim\penalty 10000\ Y_{k}\leq\dim\penalty 10000\ \penalty 10000\ \tilde{\pi}^{-1}(Z_{k})=\dim\penalty 10000\ \penalty 10000\ Z_{k}+\dim\penalty 10000\ \penalty 10000\ KGL_{n}.$
@@
$\begin{split}\operatorname{codim}_{Y}(Y_{k})=&\dim\penalty 10000\ \penalty 10000\ Y-\dim\penalty 10000\ \penalty 10000\ Y_{k}\\ &\geq\penalty 10000\ \penalty 10000\ \dim\penalty 10000\ \penalty 10000\ Y-\dim\penalty 10000\ \penalty 10000\ Z_{k}-\dim\penalty 10000\ \penalty 10000\ KGL_{n}=\dim\penalty 10000\ \penalty 10000\ Z-\dim\penalty 10000\ \penalty 10000\ Z_{k}\\ &=\operatorname{codim}_{Z}(Z_{k})>k.\end{split}$
Therefore, $\widetilde{N^{cl}_{Gie,0}}$ is an irreducible smooth stack, almost very good and $N^{cl}_{Gie,0}$ is an irreducible, almost very good stack. ∎
Theorem A.12.
The stack $M^{cl}_{Gie,0}$ is an irreducible local complete intersection of pure dimension $2\dim N_{Gie,0}+1$.